**MUST SHOW GRAPHICAL METHOD FOR CALCULATING GRADIENT**

Question: The equation for the instantaneous voltage across a discharging capacitor is given by v= Voe^-t/t, where Vo is the initial voltage and t is the time constant of the circuit..

A) draw a graph that includes a graphical method for calculating gradient of voltage against time for Vo = 12V and t= 2s, between t = 0s and t= 10s?

B) calculate the gradient at t = 2s and t= 4s? (must be done on a numbered graph)

C) differentiate v= 12e^-t/2 and calculate the value of dv/dt at t=2s and t=4s?

D) compare results for part b and c?

E) calculate the second derivative of the instantaneous voltage (d^2 v / dt^2)?
1
Expert's answer
2019-11-25T11:51:08-0500

A)




We drew a tangent at "t=2" and found the slope by finding x intercept and y intercept of the tangent

B)From the graph,

At "t=2s"

"x -intercept=2.6"

"y-intercept=3.2"

"Gradient = m= -tan(\theta)= -3.2/2.6=-1.23"

At "t=4s"

"x-intercept=4.7"

"y-intercept= 0.3"

"Gradient = m= -tan(\theta)= -0.3/4.7=-0.06"

C)"V=12e^{-t}\/t"

"dV\/dt=-12e^{-t}(1+t)\/t^2"

At "t=2s"

"dV\/dt=-1.23"

At "t=4s"

"dV\/dt=-0.07"

D)"(m_2)_c>(m_2)_b"

"(m_4)_c>(m_4)_b"

D)"dV\/dt=-12e^{-t}(1+t)\/t^2"

"d^2V\/dt^2=V_oe^{-t}(t^2+2t+2)\/t^3"